Dewey Corp. is expected to have an EBIT of $2,900,000 next year. Depreciation, the increase in net working capital, and capital spending are expected to be $225,000, $130,000, and $230,000, respectively. All are expected to grow at 17 percent per year for four years. The company currently has $17,500,000 in debt and 840,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 2.9 percent indefinitely. The company’s WACC is 9.3 percent and the tax rate is 24 percent. What is the price per share of the company's stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Expected FCF, FCF1 = EBIT * (1 - Tax Rate) + Depreciation - Increase in NWC – Capital Spending
Expected FCF, FCF1 = $2,900,000 * (1 - 0.24) + $225,000 - $130,000 - $230,000
Expected FCF, FCF1 = $2,069,000

Growth rate for next 4 years is 17% and a constant growth rate (g) of 2.90% thereafter

FCF2 = $2,069,000 * 1.170 = $2,420,730
FCF3 = $2,420,730 * 1.170 = $2,832,254
FCF4 = $2,832,254 * 1.170 = $3,313,737
FCF5 = $3,313,737 * 1.170 = $3,877,072
FCF6 = $3,877,072 * 1.029 = $3,989,507

WACC = 9.30%

Horizon Value = FCF6 / (WACC - g)
Horizon Value = $3,989,507 / (0.0930 - 0.0290)
Horizon Value = $62,336,047

Value of Firm = $2,069,000/1.093 + $2,420,730/1.093^2 + $2,832,254/1.093^3 + $3,313,737/1.093^4 + $3,877,072/1.093^5 + $62,336,047/1.093^5
Value of Firm = $50,856,825.845

Value of Equity = Value of Firm - Value of Debt
Value of Equity = $50,856,825.845 - $17,500,000
Value of Equity = $33,356,825.845

Stock Price = Value of Equity / Number of Shares Outstanding
Stock Price = $33,356,825.845 / 840,000
Stock Price = $39.71